# Faster Neighborhood Attention: Reducing the $\mathcal{O}(n^{2})$ Cost of Self Attention at the Threadblock Level

### v2

@@ Abstract
Neighborhood attention [8] reduces the cost of self attention by restricting each token’s attention span to its nearest neighbors. This restriction, parameterized by a window size and dilation factor, draws a spectrum of possible attention patterns between linear projection and self attention. Neighborhood attention, and more generally sliding window attention patterns, have long been bounded by infrastructure, particularly in higher-rank spaces (2-D and 3-D), calling for the development of custom kernels, which have been limited in either functionality, or performance, if not both. In this work, we aim to massively improve upon existing infrastructure by providing two new methods for implementing neighborhood attention. We first show that neighborhood attention can be represented as a batched GEMM problem, similar to standard attention, and implement it for 1-D and 2-D neighborhood attention. These kernels on average provide 895% and 272% improvement in full precision latency compared to existing naive CUDA kernels for 1-D and 2-D neighborhood attention respectively. We find certain inherent inefficiencies in all unfused implementations of neighborhood attention that bound their performance and hinder their lower-precision scalability. Motivated by the progress made into fused dot-product attention kernels, we developed fused neighborhood attention; an adaptation of fused dot-product attention kernels that allow fine-grained control over attention across different spatial axes. Known for reducing the quadratic time complexity of self attention to a linear complexity, neighborhood attention can now enjoy a reduced and constant memory footprint, and record-breaking half precision latency. We observe that our fused implementation successfully circumvents some of the unavoidable inefficiencies in unfused implementations. While our unfused GEMM-based kernels only improve half precision performance compared to naive kernels by an average of 496% and 113% in 1-D and 2-D problems respectively, our fused kernels improve naive kernels by an average of 1607% and 581% in 1-D and 2-D problems respectively. These translate into up to 97% improvement in inference speed in existing models based on neighborhood attention, and additionally extend its applicability in vision (image and video) as well as other modalities. We will open source our kernels at github.com/SHI-Labs/NATTEN.
@@
Figure 1: Overview of average improvement in speed on Ampere from our proposed implementation. Baseline is the set of naive CUDA kernels introduced in Neighborhood Attention Transformer [8]. GEMM-based NA improves 1-D problems by an average of 895% (FP32) and 496% (FP16), and 2-D problems by an average of 272% (FP32) and 113% (FP16). GEMM-based NA does not implement 3-D problems yet. Fused NA boosts performance further and improves 1-D problems by an average of 1072% (FP32) and 1607% (FP16), 2-D problems by an average of 381% (FP32) and 581% (FP16), and 3-D problems by an average of 460% (FP32) and 552% (FP16).
@@ 1 Introduction
Inarguably among the most highly utilized and influential primitives in modern deep learning, attention has long been cited for its complexity and memory footprint, especially when the query and context sets are identical (self attention). For years since its adoption in deep learning [20], the most common implementation of attention was through two batched GEMM (General Matrix-Matrix Multiplication) operations, sometimes referred to as “BMM-style” attention. This implementation stores attention weights to global memory, which can become a bottleneck in both speed and memory footprint. As the number of possible attention weights grow, the problem gets bounded by global memory bandwidth, as well as the amount of global memory required.
@@
Over the past few years, some works proposed attention implementations in which attention weights are not stored to global memory, but instead kept in on-chip memory (shared memory or register file) [16, 5], which are then directly applied to corresponding values, writing the resulting attention outputs to global memory. These implementations, known as fused or memory-efficient attention, reduce the number of global memory accesses in addition to global memory usage. Thanks to the first open-source implementation, Flash Attention [5], these fused attention kernels have started replacing the standard BMM-style implementations in many deep learning frameworks and inference engines such as PyTorch [14].
@@
Orthogonal to these efforts, many have sought to address the quadratic complexity in self attention, which is the more predominantly used operation in vision models. Neighborhood attention [8] is one such method in which each query token is restricted to only interact with its nearest neighboring context tokens. In most cases, this pattern creates a sliding window pattern, which in deep learning can also be seen in the convolution operator. This restriction can similarly be parameterized by a window size and dilation factor, and reduces the quadratic complexity of self attention down to a linear complexity. This reduction is however very difficult to implement at the tensor library or deep learning framework level. Tensor views can represent sliding window attention [17], but not the neighborhood attention pattern. In addition, GEMM operators do not typically support taking in the sliding window views in higher-rank spaces (2-D and 3-D) without explicit copying, which in practice undoes the theoretical efficiency gain from the reduced attention complexity. As a result, neighborhood attention was proposed along with an extension carrying naive CUDA kernels [8] implementing the two BMM operators. While those kernels can provide competitive FP32 performance, and in some cases even FP16/BF16 performance, they fall short of being generally adoptable in larger scale experiments. In addition, fused attention kernels such as Flash Attention effectively eliminate the $\mathcal{O}(n^{2})$ memory footprint in self attention, while also reducing latency significantly [5], making subquadratic attention patterns that are only possible to implement in “BMM-style” less feasible in practice.
@@
In this work, we present two new classes of neighborhood attention kernels: GEMM-based BMM-style kernels (GEMM NA), and fused kernels (Fused NA), which are aimed at providing significantly improved infrastructure for neighborhood attention. We first show that neighborhood attention, and by extension sliding window attention, both of which are GEMV (General Matrix-Vector Multiplication) problems, can be expressed as GEMM problems with space-aware tiling and gather/scatter fusion. This would allow implementing such attention patterns with GEMMs, which in turn can benefit greatly from the underlying efficiency of GEMM kernels. We then extend the same logic to fused attention kernels by factoring out all logic that corresponds to the “token space”, and writing specializations that assume higher rank spaces, such as 2-D and 3-D. This, in theory, allows any fused attention kernel to be modified to accommodate token spaces of any rank. In addition, most of the logic is evaluated at compile time, resulting in minimal overhead stemming from overly complicated kernels, branch divergence, or any of the like. Finally, the structural simplicity of the resulting fused neighborhood attention kernels allows for easily adding features such as varying window size / dilation values across spatial axes, causal masking, and more.
@@
Figure 2: Illustration of the spectrum of possible attention patterns provided by neighborhood attention. Neighborhood attention attempts to center the query token (red) within the context window (blue), as opposed to standard sliding window attention, which forces it. Neighborhood attention with window size 1 is equivalent to a linear projection (no attention). Neighborhood attention approaches self attention as window size grows, and matches it when equal to input size. Dilated windows can introduce more global context without any additional compute. Causal neighborhood attention shifts the neighborhood to end with the query.
@@ 2 Related works
Attention being adopted as a primitive is largely owed to the Transformer architecture [20], which despite its original application in machine translation rose to a predominant deep learning architecture, in addition to its design and use of the attention operator extending across applications and modalities [13, 6, 1, 15]. Attention is defined as an operation between a query set and a context set, both of which undergo linear projections, with the latter projected into a pair of key and value sets. Scaled dot product of queries and keys, $A$, is mapped into a probability distribution through the softmax operator, which produces the final attention weights, $P$. The output is a set of tokens, each derived from the weighted sum of all value tokens according to the token’s attention weights. It can be formally expressed as:
@@
where $\sqrt{d}$ is the scale term, and $d$ is the number of dimensions for which the dot product is computed (the embedding dimension of $Q$ and $K$). Dot product self attention, or simply, self attention, is a special case in which the query and context sets are identical. This means for a set of $n$ tokens, the attention weights matrix, $P$, is $\in\mathbb{R}^{n\times n}$, incurring an $\mathcal{O}(n^{2})$ time and space complexity. In addition, the softmax term requires reducing entire rows of the attention weight matrix, making kernel fusion much more challenging. However, fused attention kernels would at the very least eliminate the $\mathcal{O}(n^{2})$ global memory footprint, itself a significant achievement that would pave the way for the application of attention across modalities. To our knowledge, the first open-source implementation of a fused multi-headed attention (FMHA) kernel was contributed to the NVIDIA Apex  [^1] https://github.com/NVIDIA/apex project by Young-Jun Ko, which was primarily used for accelerating certain Transformer-based language models, but heavily limited in terms of supported problem sizes. Milakov and Gimelshein [12] presented a technique for computing partial softmax statistics, which when aggregation over all subsets of the original set results in exact softmax computation. This method makes the fusion of attention kernels more practical, because they would no longer be required to compute a full row of attention weights [16]. Dao et al. [5] presented and open-sourced Flash Attention, a generic fused attention kernel along with an efficient backward-pass kernel, both of which outperformed standard BMM-style attention implementations available. It was quickly adopted by many standard libraries, such as PyTorch [14], and further improved in the more recent Flash Attention V2 [4].
@@
Parallel to these efforts, many proposed restricted attention patterns, in which the context for every query token is restricted in order to generate fewer attention weights, which in turn reduces the $\mathcal{O}(n^{2})$ time and space complexity. Stand-alone self attention (SASA) [17] is a simple 2-dimensional sliding window attention pattern, which was shown to effectively replace convolution operations in ResNet [9] variants. Noting challenges in implementing such patterns without incurring additional overhead from tensor copies and expansion, the authors later moved away from explicit sliding window attention patterns to alternatives that relaxed the sliding window movement in HaloNet [19]. In addition to these works, sliding window attention patterns in 1-dimensional spaces has been explored in language, in works such as Sparse Transformers [3], Longformer [2], BigBird [22], and more recently, Mistral [10]. Neighborhood attention [8, 7] is the practice of restricting the context of each token to its nearest neighbors, which in many cases behaves like a sliding window pattern, with the exception of corner cases in which the query cannot be centered in a sliding window. Per the original definitions from SASA [17], and Longformer [2], the sliding window can go out of bounds, in which case the attention weights corresponding to out-of-bounds context tokens are masked. This means tokens close to spatial bounds interact with fewer context tokens. This difference allows neighborhood attention to approach self attention as window size grows. In addition, neighborhood attention defines a dilation factor [7], where the number of such corner cases only increase. Fig. 2 depicts possible attention pattern for a single token under different neighborhood attention parameters. Facing similar implementation challenges as previous works, particularly SASA [17], neighborhood attention was implemented with naive CUDA kernels, and packaged as a PyTorch extension, $\mathcal{N}ATTEN$. While those kernels have accelerated research in this direction, they simply were not intended to fully utilize the underlying hardware. The only exception is the tiled kernels, which are somewhat better optimized, but only implement the $QK^{T}$ operation for 2-dimensional problems with $32$ dimensions per attention head, and window sizes up to $13\times 13$. In addition, with the rise of fused attention kernels such as Flash Attention [5], such implementations would have a hard time competing in terms of latency, and in most cases unable to compete in terms of global memory footprint.
@@
To address these challenges, we present two new implementations and integrate them into $\mathcal{N}ATTEN$, aiming to accelerate all neighborhood attention applications, reduce their existing memory overhead, and extend existing functionality. We first simplify the operations that implement neighborhood attention’s forward and backward pass into 3 primary operators, and show each can be implemented with a batched GEMM kernel with either a fused gather or scatter operation. We then point out key limitations in unfused neighborhood attention implementations that would prevent them from achieving competitive performance compared to standard BMM-style attention implementations. Motivated by this, and the progress made in fused attention kernels, we propose fused neighborhood attention, borrowing from some of the data loading and tiling logic in our batched GEMM methodology. We implement both approaches using CUTLASS [18], and show that the batched GEMM kernels can successfully outperform most existing $\mathcal{N}ATTEN$ kernels in performance, and that our fused kernels can outperform our batched GEMM kernels while reducing the memory footprint.
@@ 3 Methodology
Figure 3: A simplified illustration of our of GEMM-based implementation of the 2-D $QK^{T}$ operation. Queries are tiled according to their original spatial rank, which in this case is 2. Keys are first restricted according to the context window size, which is just a haloed query tile. Depending on the GEMM configuration, the haloed key tile may or may not be further tiled into smaller tiles compatible with the underlying GEMM routine. Once the GEMM computes the dot product between the query and key tiles, the epilogue will scatter valid attention weights into the attention weights tensor, and discards the rest.
Herein we describe the 3 primary operations (excluding softmax) that are required to implement a full neighborhood attention forward and backward pass. We then show that each operation can be expressed as a batched GEMM problem, as long as tiling is done according to the underlying spatial rank, and attention weights are scattered and gathered. The latter is a major bottleneck for all BMM-style implementations of neighborhood attention, which in particular limits their scalability on modern hardware (i.e. Ampere and Hopper architectures.) We then introduce our fused neighborhood attention formulation which builds upon our batched GEMM formulation and tiles according to the underlying spatial rank. This approach no longer requires gathering or scattering attention weights to global memory by definition, and thereby circumvents the aforementioned bottleneck and boosts neighborhood attention performance on modern architectures.

### v3

@@ Abstract
Neighborhood attention reduces the cost of self attention by restricting each token’s attention span to its nearest neighbors. This restriction, parameterized by a window size and dilation factor, draws a spectrum of possible attention patterns between linear projection and self attention. Neighborhood attention, and more generally sliding window attention patterns, have long been bounded by infrastructure, particularly in higher-rank spaces (2-D and 3-D), calling for the development of custom kernels, which have been limited in either functionality, or performance, if not both. In this work, we aim to massively improve upon existing infrastructure by providing two new methods for implementing neighborhood attention. We first show that neighborhood attention can be represented as a batched GEMM problem, similar to standard attention, and implement it for 1-D and 2-D neighborhood attention. These kernels on average provide 895% and 272% improvement in full precision runtime compared to existing naive CUDA kernels for 1-D and 2-D neighborhood attention respectively. We find that aside from being heavily bound by memory bandwidth, certain inherent inefficiencies exist in all unfused implementations of neighborhood attention, which in most cases undo their theoretical efficiency gain. Motivated by the progress made into fused dot-product attention kernels, we developed fused neighborhood attention; an adaptation of fused dot-product attention kernels that allow fine-grained control over attention across different spatial axes. Known for reducing the quadratic time complexity of self attention to a linear complexity, neighborhood attention can now enjoy a reduced and constant memory footprint, and record-breaking half precision runtime. We observe that our fused implementation successfully circumvents some of the unavoidable inefficiencies in unfused implementations. While our unfused GEMM-based kernels only improve half precision performance compared to naive kernels by an average of 548% and 193% in 1-D and 2-D problems respectively, our fused kernels improve naive kernels by an average of 1759% and 958% in 1-D and 2-D problems respectively. These improvements translate into up to 104% improvement in inference and 39% improvement in training existing models based on neighborhood attention, and additionally extend its applicability to image and video perception, as well as other modalities. Our work is open-sourced at https://github.com/SHI-Labs/NATTEN/.
@@
Figure 1: Overview of average improvement in speed on A100 from our proposed implementation. Baseline is the set of naive CUDA kernels introduced in Neighborhood Attention Transformer [9]. GEMM-based NA improves 1-D problems by an average of 548% (forward pass) and 502% (forward + backward), and 2-D problems by an average of 193% (forward pass) and 92% (forward + backward). GEMM-based NA does not implement 3-D problems yet. Fused NA boosts performance further and improves 1-D problems by an average of 1759% (forward pass) and 844% (forward + backward), and 2-D problems by an average of 958% (forward pass) and 385% (forward + backward), and 3-D problems by an average of 1135% (forward pass) and 447% (forward + backward).
@@ 1 Introduction
Inarguably among the most highly utilized and influential primitives in modern deep learning, attention has long been cited for its complexity and memory footprint, especially when the query and context sets are identical (self attention). For years since its adoption in deep learning [23], the most common implementation of attention was through two batched GEMM (General Matrix-Matrix Multiplication) operations, sometimes referred to as “BMM-style” attention. This implementation stores attention weights to global memory, which can become a bottleneck in both speed and memory footprint. As the number of tokens grow, the number of attention weights grow as well, and the problem gets bounded by global memory bandwidth and capacity.
@@
Over the past few years, some works proposed attention implementations in which attention weights are kept in on-chip memory (shared memory or register file) instead, until the second matrix multiplication is performed and the resulting attention outputs are written directly to global memory  [18, 6]. These implementations, known as fused or memory-efficient attention, reduce the number of global memory accesses in addition to global memory usage, and successfuly turn dot product attention into a compute-bound problem at scale. Thanks to the first open-source implementation, Flash Attention [6], these fused attention kernels have started replacing the standard BMM-style implementations in many deep learning frameworks and inference engines such as PyTorch [16].
@@
Orthogonal to these efforts, many have sought to address the quadratic complexity of self attention, which can become a significant bottleneck in vision models more quickly. Neighborhood attention [9] is one such method in which each query token is restricted to only interact with its nearest neighboring context tokens. In most cases, this pattern creates a sliding window pattern, like that of the discrete convolution operator heavily employed in vision models. This restriction can similarly be parameterized by a window size and dilation factor, and reduces the quadratic complexity of self attention down to a linear complexity. This approach is, however, very difficult to implement at the tensor library or deep learning framework level. Tensor views can represent sliding window attention [19], but not the neighborhood attention pattern. In addition, standard GEMM implementations typically do not support such tensor views in higher-rank/multi-dimensional spaces (2-D and 3-D) without explicit copying into contiguous tensors, which in practice undoes the theoretical efficiency gain from the reduced attention complexity. As a result, neighborhood attention was proposed along with an extension carrying naive CUDA kernels [9] implementing the operation. While those kernels provided competitive FP32 performance in eager mode inference, and in some cases even FP16/BF16 performance, they fall short of general adoption in larger scale experiments. In addition, fused attention implementations, such as Flash Attention, effectively eliminate the $\mathcal{O}(n^{2})$ memory footprint in self attention, while also reducing runtime significantly [6], making subquadratic attention patterns that are only possible to implement “BMM-style” less practical.
@@
In this work, we present two new classes of neighborhood attention kernels: GEMM-based BMM-style kernels (GEMM NA), and fused kernels (Fused NA), which are aimed at providing significantly improved infrastructure for neighborhood attention. We first show that neighborhood attention, and by extension sliding window attention, both of which are GEMV (General Matrix-Vector Multiplication) problems, can be expressed as GEMM problems with space-aware tiling and gather/scatter fusion. This would allow implementing such attention patterns with performance-optimized GEMM primitives, which can also utilize specialized hardware components such as NVIDIA’s Tensor Cores. We then extend the same logic to fused attention kernels by removing all assumptions that the token mode (“space”) is rank-1 (single-axis). We write specializations that support higher-rank/multi-dimensional spaces, such as 2-D and 3-D. This, in theory, allows any fused attention kernel to be modified to accommodate token spaces of any rank. In addition, part of the logic is evaluated at compile time, resulting in less overhead. Finally, the structural simplicity of the resulting fused neighborhood attention kernels allows for easily adding features such as varying window sizes / dilation values across ranks/axes, causal masking, and more.
@@
Figure 2: Illustration of the spectrum of possible attention patterns provided by neighborhood attention. Neighborhood attention only attempts to center the query token (red) within the context window (blue), unlike sliding window attention [19] which forces it. Neighborhood attention with window size 1 is equivalent to linear projection (“no attention”). Neighborhood attention approaches self attention as window size grows, and matches it when equal to input size. Dilation introduces sparse global context, and causal masking prevents interaction between query tokens that have a smaller coordinate than neighboring context tokens along the corresponding mode. Window size, dilation, and whether or not causally masked, can be defined per mode/axis.
@@ 2 Related works
Attention being adopted as a deep learning primitive is largely owed to the Transformer architecture [23], which despite its original application in machine translation rose to the position of being the predominant deep learning architecture. Its design and use of the attention operator have been extended to many other applications and modalities [15, 7, 1, 17]. Attention is defined as an operation between a two sets of vectors: a query set and a context set. The two undergo linear projections, with the latter projected into a set of key and value pairs. Scaled dot product of query and key vectors, $A$, is mapped into a probability distribution through the softmax operator, which produces the final attention weights, $P$. The output is a set of vectors, each derived from the weighted sum of all value vectors according to the query vector’s attention weights. It can be expressed as follows:
@@
where $Q$, $K$, and $V$ are matrices of query, key, and value vectors as rows respectively, $\sqrt{d}$ is the scale term, and $d$ is the number of dimensions for which the dot product is computed (number of columns in $Q$ and $K$). Dot product self attention, or simply, self attention, is a special case in which the query and context sets are identical. This means for a set of $n$ input vectors, the attention weights matrix, $P$, is $\in\mathbb{R}^{n\times n}$, incurring an $\mathcal{O}(n^{2})$ time and space complexity. In addition, the softmax term requires column-wise reduction over the attention weight matrix, making kernel fusion more challenging.
@@
Nevertheless, fused attention kernels successfully eliminate the $\mathcal{O}(n^{2})$ global memory footprint, which makes self attention finally bound by compute and not memory bandwidth. These two achievements paved the way for the scaling and application of attention across modalities. To our knowledge, the first open-source implementation of a fused multi-headed attention (FMHA) kernel was contributed to the NVIDIA Apex  [^1] https://github.com/NVIDIA/apex project by Young-Jun Ko, which was primarily used for accelerating inference of Transformer-based language models. As a result of that, it was heavily limited in terms of supported models and problem sizes, as it was performing a full softmax reduction step within the kernel. On the other hand, Milakov and Gimelshein [14] presented a technique for computing partial softmax statistics, over which we can perform a final reduction step and derive exact softmax results. This method makes the fusion of attention kernels more practical, because they would no longer be required to compute a full row of attention weights before proceeding to perform the second matrix multiplication. Dao et al. [6] presented and open-sourced Flash Attention, which utilizes online softmax in order to create a performant and generic fused attention implementation. Outperforming BMM-style implementations available in both training and inference, Flash Attention was quickly adopted by many frameworks such as PyTorch [16], and further improved for the NVIDIA Ampere [5] and Hopper architectures [20].
@@
Parallel to these efforts, many proposed restricted self attention patterns, in which context is restricted to a subset in order to generate fewer attention weights, which in turn reduces the $\mathcal{O}(n^{2})$ time and space complexity. Stand-alone self attention (SASA) [19] is a simple 2-dimensional sliding window attention pattern, which was shown to effectively replace convolution operations in ResNet [10] variants. Noting challenges in implementing such patterns without incurring additional overhead from tensor copies and expansion, the authors later moved away from explicit sliding window attention patterns to alternatives that relaxed the sliding window movement in HaloNet [22]. In addition to these works, sliding window attention patterns in 1-dimensional spaces has been explored in language, in works such as Sparse Transformers [4], Longformer [2], BigBird [26], and more recently, Mistral [11]. Neighborhood attention [9, 8] is the practice of restricting the context of each token to its nearest neighbors, which in many cases behaves like a sliding window pattern, with the exception of corner cases in which the query cannot be centered in a sliding window. Per definitions from SASA [19] and Longformer [2], the sliding context window can go out of bounds, in which case the attention weights corresponding to out-of-bounds tokens are masked. This means tokens close to spatial bounds interact with fewer context tokens. This difference allows neighborhood attention to approach self attention as window size grows. In addition, neighborhood attention defines a dilation factor [8], where the number of such corner cases only increase. Fig. 2 depicts possible attention patterns for a single token under different neighborhood attention parameters. Facing similar implementation challenges as previous works [19], neighborhood attention was implemented with naive CUDA kernels packaged as a PyTorch extension, named $\mathcal{N}ATTEN$. While those kernels have accelerated research in this direction, they were simply not intended to fully utilize the underlying hardware. The only exception is the tiled kernels, which are somewhat better optimized, but only apply to a fraction of common use cases, and are not extensible. In addition, with the rise of fused attention kernels such as Flash Attention [6], such implementations which are not performance-optimized and heavily memory-bandwidth-bound, can hardly compete in terms of performance and memory footprint.
@@
To address these challenges, we present two new implementations and integrate them into $\mathcal{N}ATTEN$, aiming to accelerate all neighborhood attention applications, reduce their existing memory overhead, and extend existing functionality. We first simplify the operations that implement neighborhood attention’s forward and backward pass into 3 primary operators, and show each can be implemented as batched GEMM kernels with a fused gather/scatter operation. We then point out key limitations in unfused neighborhood attention implementations that would prevent them from achieving competitive performance compared to standard BMM-style attention implementations (in more memory-bandwidth-bound cases.) Motivated by this, and the progress made in fused attention kernels, we propose fused neighborhood attention, which directly extends our batched GEMM methodology. Since our main objectives are efficiency, Tensor Core utilization, and performance optimization, and our methodology requires significant flexibility in the programming model, we implement both approaches in CUDA C++ using NVIDIA’s CUTLASS [21] framework. We show that the batched GEMM kernels can successfully outperform most existing $\mathcal{N}ATTEN$ kernels in performance, and that our fused kernels can outperform our batched GEMM kernels while reducing the memory footprint.
@@ 3 Methodology
Figure 3: Illustration of our GEMM-based implementation of the 2-D PN operation. Input tensors $Q$ and $K$ are tiled according to their 2-D spatial layout. $Q$ is tiled with a static tile shape, $T_{h}\times T_{w}$. $K$ is tiled with a haloed shape of the $Q$ tile, $T^{\prime}_{h}\times T^{\prime}_{w}$, which is a function of the attention window size ($k_{h}\times k_{w}$) and the $Q$ tile coordinates. Once tiles are moved into local memory, they are viewed in matrix layout, and a $T_{h}T_{w}\times T^{\prime}_{h}T^{\prime}_{w}\times d$ shaped GEMM is computed ($d$ is embedding dim). Once done, the tile of dot products with shape $T_{h}T_{w}\times T^{\prime}_{h}T^{\prime}_{w}$ is scattered into valid attention weights of shape $T_{h}\times T_{w}\times k_{h}k_{w}$.
Herein we describe three primary operations (excluding softmax) that are required to implement a full neighborhood attention forward and backward pass. We then show that each operation can be expressed as a batched GEMM problem, as long as tiling is done according to the underlying spatial rank, and attention weights are scatter/gathered. However, we find that scatter/gather is a major bottleneck for all unfused implementations of neighborhood attention, limiting their low-precision performance specifically on more recent architectures (Ampere and later.) We then introduce our fused neighborhood attention (FNA) formulation, which builds on our batched GEMM formulation and tiles according to the underlying spatial rank. This approach no longer requires scatter/gathering of attention weights to/from global memory by definition, and thereby circumvents the aforementioned bottleneck and successfully boosts lower-precision performance on modern architectures.